OpenStudy (anonymous):

its a integration question find the volume of ellipsoid by rotating the ellipse x^2/a^2 + y^2/b^2=1

OpenStudy (anonymous):

\[x^2/a^2+ y^2/b^2 =1\]

OpenStudy (anonymous):

On rotating the ellipse around the x-axis, an ellipsoid is formed. Consider a thin cylinder of thickness dx at a distance of x from the origin whose radius = y Volume of this thin cylinder, dV = πy^2 dx => Volume of the ellipsoid, V = ∫ (x = - a to x = a) dV = ∫ (x = - a to x = a) πy^2 dx = ∫ (x = - a to x = a) π * (b^2/a^2) (a^2 - x^2) dx = (πb^2/a^2) [a^2x - x^3/3] ... [x = - a to x = a] = (πb^2/a^2) [(a^3 - a^3/3) - (- a^3 + a^3/3)] = (πb^2/a^2) * (4/3) a^3 = (4/3) πab^2.

OpenStudy (anonymous):

whoa thanks you solved the whole problem :D

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

Man that is painful on my eyes. I need latex.

OpenStudy (opcode):

Since I take it you're all new. You can make mathematical text look nice using \(LaTeX...\) \(\large \frac{x^2}{a^2} + \frac {y^2}{b^2}= 1\) Code for that: ``` \(\large \frac{x^2}{a^2} + \frac {y^2}{b^2}= 1\) ``` So basically if you want math text to look nice input it in between. \(\(\) and \)

OpenStudy (anonymous):

How do you add code!!?

OpenStudy (anonymous):

help me with the diagram though :|

hartnn (hartnn):

add code using ` like \(` aa `\) will give you `aa`

OpenStudy (anonymous):

For inline math use `\(` and `\)`. For full math use `\[` and `\]`

OpenStudy (anonymous):

-_-

hartnn (hartnn):

we are deviating from the question!! someone help him with the diagram!

OpenStudy (anonymous):

are you suree???

OpenStudy (anonymous):

there is no area bounded by them,so volume either

OpenStudy (anonymous):

` | へ  へ  ` ` | の の \\ ` ` | も ` ` | へ | ` ` \__________/ `

OpenStudy (anonymous):

cute

OpenStudy (opcode):

On rotating the ellipse around the x-axis, an ellipsoid is formed. Consider a thin cylinder of thickness dx at a distance of x from the origin whose radius = y Volume of this thin cylinder, \(dV = \pi y^2 dx\) => Volume of the ellipsoid, V \(\Large= ∫ (x = - a\) to \(x = a) dV\) \(\Large= ∫ (x = - a\) to x = a) \(\pi y^2 dx\) \(\Large= ∫ (x = - a\) to \(x = a) \pi * (\frac{b^2}{a^2}) (a^2 - x^2) dx\) \(\Large= (\frac{\pi b^2}{a^2}) [a^2x - \frac{x^3}{3}]\) ... \([x = - a\) to \(x = a]\) \(\Large= (\frac{\pi b^2}{a^2}) [(a^3 - \frac{a^3}{3}) - (- a^3 + \frac{a^3}{3})]\) \(\Large= (\frac{\pi b^2}{a^2}) * (\frac{4}{3}) a^3\) \(\Large= (\frac{4}{3}) \pi ab^2.\) @wio \(LaTeX!\)

OpenStudy (anonymous):

∫(x=−a to x=a)dV should be \[ \int_{-a}^adV \] `\[ \int_{-a}^adV \]`

hartnn (hartnn):

how about \[ \int \limits _{-a}^adV \] `\[ \int \limits _{-a}^adV \]`

OpenStudy (anonymous):